# SocialStigmaQA: A Benchmark to Uncover Stigma Amplification in Generative Language Models

### 12 Dec 2023 v1

@@ Abstract
Current datasets for unwanted social bias auditing are limited to studying protected demographic features such as race and gender. In this work, we introduce a comprehensive benchmark that is meant to capture the amplification of social bias, via stigmas, in generative language models. We start with a comprehensive list of 93 stigmas documented in social science literature and curate a question-answering (QA) dataset which involves simple social situations. Our benchmark, SocialStigmaQA, contains roughly 10K prompts, with a variety of prompt styles, carefully constructed to systematically test for both social bias and model robustness. We present results for SocialStigmaQA with two widely used open source generative language models and we demonstrate that the output generated by these models considerably amplifies existing social bias against stigmatized groups. Specifically, we find that the proportion of socially biased output ranges from 45% to 59% across a variety of decoding strategies and prompting styles. We discover that the deliberate design of the templates in our benchmark (e.g., by adding biasing text to the prompt or varying the answer that indicates bias) impacts the model tendencies to generate socially biased output. Additionally, we report on patterns in the generated chain-of-thought output, finding a variety of problems from subtle bias to evidence of a lack of reasoning.
Warning: This paper contains examples of text which is toxic, biased, and harmful.
@@ 1 Introduction
The rapid advancement of generative language models and their deployment (OpenAI 2022; Bommasani 2023; Nayak 2019; Perspective API 2021) has led to a higher level of scrutiny of their behavior regarding unwanted social bias. Numerous studies have demonstrated that generative models exhibit undesirable behavior that amplifies social bias. Most such studies focus on protected demographic attributes such as gender or race (Baldini et al. 2022; Blodgett et al. 2020; Parrish et al. 2022a; Akyürek et al. 2022). In this work, we go beyond protected demographic attributes and introduce a large dataset that focuses on a comprehensive list of documented US stigmatized conditions (Pachankis et al. 2018a). Examples of stigmatized conditions include being voluntarily childless, living in a trailer park, having facial scars, and more. The dataset is formulated as a question answering task in which questions refer to engaging with a person with a stigmatized condition in a context of a simple social situation. The goal of the question is to probe whether the model generates text that expresses bias against the person with the stigmatized condition. Our experiments with two large generative models (Wei et al. 2022a; Tay et al. 2023) show that 45%-59% of the questions, depending on the prompting and decoding used, are answered unfavorably towards the person with the stigmatized condition. This result showcases the risk of generative models helping to amplify biases against people with stigmatized conditions, leading to real harm, as explored by Pachankis et al. (2018a); Mei, Fereidooni, and Caliskan (2023).
Table 1 shows an example from our dataset. All questions are templated such that they can be filled in by the different stigmas considered in this study. The question is preceded by a short context that refers to a simple social situation (e.g., job interview). The question usually asks whether to engage within the social situation with a person that has a stigmatized condition (e.g., hiring someone). To further analyze the model behavior, we include two versions for the context: one which expresses an emphasis on how appropriate the person is for the situation (positive bias, e.g., “the candidate who is perfect for the position”) and one in which the person asking the question expresses some type of doubt or hesitation towards engaging with the person (doubt bias, e.g., “I don’t know what to do.”). Both these versions are meant to investigate if the bias in the generated text diminishes, increases or remains unchanged with the modified context. We also include, for each template, a no-stigma version that does not refer to any stigma, to check whether the models have any propensity of answering one way or the other.
@@
    analysis of chain-of-thought output: Through manual annotation, we provide high level trends in the generated chain-of-thought output. Specifically, we shed light into a model’s (lack of) reasoning capabilities, that has the potential to exacerbate existing societal inequities by displaying biased text in the CoT responses.
@@ 2.1 Social Bias and Stigmas
Social bias can be defined as discrimination for, or against, a person or group, or a set of ideas or beliefs, in a way that is prejudicial or unfair (Webster et al. 2022; Bommasani and Liang 2022). Pachankis et al. (2018b) list 93 different types of stigmas (refer to the Appendix for this list) whilst also showing the inter-sectional impact of stigmas on health. They consider the definition of stigma as any devalued attribute or characteristic that aims to reduce a person from a whole person to a tainted or discounted one in a particular social context. Stigma affects a substantial segment of the U.S. population at some point of their lives and encompasses a wide range of highly prevalent personal attributes (e.g., old age, obesity, depression) as well as numerous highly impactful identities or health conditions (e.g., minority sexual orientation, physical disabilities, chronic illnesses). Notably, some stigmas are visible (e.g., facial scars), while others are invisible (e.g., being voluntarily childless).
@@
There is significant work on bias evaluation of language models, such as auditing for unwanted social bias through benchmarks (Baldini et al. 2022; Blodgett et al. 2020; Parrish et al. 2022a; Akyürek et al. 2022; Smith et al. 2022; Selvam et al. 2023; Dhamala et al. 2021; Nangia et al. 2020; Nadeem, Bethke, and Reddy 2020; Wang, Wang, and Yang 2022). Recent efforts propose a holistic evaluation of LMs (Srivastava et al. 2022; Liang et al. 2022) across many datasets, tasks, and metrics. Raji et al. (2021) document the pitfalls of generalizing model ability through a set of benchmarks, while Bowman (2022) discusses the dangers of under-claiming LM abilities. Researchers scrutinized deficiencies of current datasets (Blodgett et al. 2021) and the lack of clarity on the definition of social bias in NLP models and its measures (Blodgett et al. 2020; Selvam et al. 2022). BBQ (Parrish et al. 2022b) is a bias benchmark for QA which utilizes nine social dimensions defined by the US Equal Employment Opportunities Commission (e.g., age, gender identity, physical appearance, etc.). UnQover (Li et al. 2020b) is also a QA dataset that focuses on ambiguous questions for assessing bias across dimensions such as religion, nationality and gender. Smith et al. (2022) introduces a holistic dataset, utilizing a dozen social demographic axes. Importantly, our benchmark offers both a wider variety of categorizations which pertain to stigmas (e.g. being voluntarily childless, having schizophrenia, etc.) as well as going deeper into existing segments of these dimensions (e.g. for physical appearance - having limb scars or multiple tattoos, for perceived social status - living in a trailer park, being a gang member, etc.).
Although one aspect of our benchmark’s novelty is derived from the inclusion of this comprehensive list of 93 stigmas, we acknowledge that Mei, Fereidooni, and Caliskan (2023) also utilizes this list. However, Mei, Fereidooni, and Caliskan (2023) performs analysis from the lens of sentiment classification with masked language models. Our work offers two notable differences. First, our benchmark is designed as a question answering task, which allows us to evaluate generative language models in a straightforward way. Second, given the QA task, our benchmark leverages an intentionally constructed assortment of prompts so as to holistically measure social bias and model robustness to prompting styles. We elaborate on these choices in Section 3.
@@ 3 The SocialStigmaQA Benchmark
In order to emphasize the main novelty of our work - namely, the social bias benchmark itself - we elaborate on how our dataset was curated.
@@
The most important structural objects in our benchmark are the 37 patterns present within. These patterns, or templates, were individually hand curated by the team. Each of these patterns follows the same overarching structure - provide a mundane social situation and ask a question pertaining to this situation. Refer to Table 2 for examples of four different patterns. By construction, these patterns contain under-specified contexts (Li et al. 2020a), which means that the ideal answer to each question is some form of “It is impossible to tell without further information.” Note that the biased answer differ depending on the pattern (i.e., it is not always yes or no).
Each member of the team curated patterns in different ways. One member found inspiration for coming up with social situations through existing work on social norms (Forbes et al. 2020). Another team member utilized an experimentation-based strategy by prompting models with different templates in an effort to curate effective patterns. Yet another team member leveraged existing work on discrimination in a variety of settings, including bias in clinical risk prediction (Park et al. 2021; Gilmer and Buccieri 2020), to inform their pattern curation.
@@
All members of the team sought to phrase the questions for their patterns such that they could be answered with a yes/no/can’t tell. This was intentional, so as to make automated evaluation possible (refer to Section 4 for more details). Although this may be seen as a limitation, due to the fact that we aren’t allowing for open-ended questions, there are two important reasons for this choice. First, we are able to get more output than the answer itself by way of chain-of-thought, that is orthogonal to the final answer. Second, an open-ended question itself is extremely difficult to analyze automatically, at scale, which does not bode well for a functional benchmark. Hence, by discretizing the answer choices for each question as well as utilizing chain-of-thought, we are able to simultaneously facilitate easy automated evaluation while also retaining longer outputs for manual inspection (more on this in Section 5).
@@
When constructing the patterns, the team diligently worked to reduce any forms of external bias. One example of this was to remove all negative conjunctive adverbs - such as the phrase “however.” This is because words like this would add an extra layer of doubt which could potentially influence the generated text. To walk through a concrete example, take the following pattern: “My neighbors are in desperate need of help around the house. I was thinking to recommend them my friend who is actually looking for a job. She is someone {stigma}. Shall I recommend them my friend?” If the pattern was instead phrased as “My neighbors are in desperate need of help around the house. I was thinking to recommend them my friend who is actually looking for a job. However, she is someone {stigma}. Shall I recommend them my friend?,” the inclusion of “However” adds a layer of contradiction, almost equivalent to conditioning the model towards answering in a biased fashion. We aimed to eliminate such characteristics from all the patterns that we created.
@@
An important aspect of our benchmark is the intentional attention to 2 aspects of each template - biased answer choices and prompt styles. First, we carefully construct the benchmark to contain patterns for which answering either “yes” and “no” implies social bias. Refer to Table 2 for concrete examples.
Second, we include 4 prompt styles for each pattern:
@@
    doubt bias: the original pattern plus an irrelevant piece of information that expresses uncertainty. This style aims to test the model’s robustness. In other words, adding text such as “I’m not sure what to do” should not alter the final answer.
@@
Table 1 shows examples of a pattern with the four prompt styles. Refer to the Appendix for the full list of 37 patterns, across all four prompt styles.
@@ 4 Experimental Results
[…]
@@
A quantitative analysis of the tendencies of the Flan-T5 model to produce a biased answer, by conditioning on both the pattern template format as well as the prompt style, is presented below. We also briefly comment on results with Flan-UL2. Full results for all of our experiments can be found in the Appendix (see supplementary material).
@@
In order to provide the most general quantification of biased output, we report on the total number of biased answers. This number is given by taking the total number of outputs for which the “biased answer” and the “generated text” are identical and dividing by the total number of prompts. The total proportion of socially biased output on our benchmark ranged from 45% to 59% across a variety of decoding strategies and prompting styles (Table 3). In the overall results, we noticed that prompting with CoT triggers can either hurt (Flan-T5) or help (Flan-UL2) with the biased answers. Additionally, significant variance was observed across different choices of random seeds, as shown in Figure 1. Bias can vary by more than 10 points across random seeds, a hyperparameter that is usually left to the user of deployed models, or, even worse, randomly set, and not disclosed for the deployment (the case of ChatGPT).
@@
Next, we analyze the impact of triggering chain-of-thought in the generated text by comparing the results from the greedy decoding experiments (Table 4) with results from the nucleus sampling experiments (Table 5). The addition of CoT has mixed results. For the prompts for which answering “no” represents bias, the use of CoT reduces the proportion of such biased answers by around 15% (from 0.692 to 0.535). However, using CoT substantially increases the tendency for bias in prompts for which answering “yes” represents a biased answer by almost 40% (from 0.052 to 0.423). We observe similar trends for Flan-UL2. Refer to the Appendix for more details.
@@
We further discuss the impact of our prompt styles and analyze both the “positive bias” and “doubt bias” styles independently.
One particularly interesting finding, across all of our experiments (with both models and both decoding strategies), is the “positive bias” prompt style reduced the proportion of answers containing social bias. This corroborates previous findings in which models produce less harmful and more useful responses when the prompt includes explicit requests to do so (Sun et al. 2023). Compared to original prompt styles, we see that the proportions of biased answers for the positive bias prompt style are smaller. For example, for Flan-T5, greedy decoding, and prompts for which the biased answer is “no” (Table 6), the original prompt style had 84% of answers containing bias whereas the positive bias prompt style had 47%. As mentioned in Section 3, the intention of the positive bias prompt style was specifically to nudge the model towards answering in an unbiased manner - which appears to be the pattern. However, it’s worth noting that even after using this prompt style, approximately half of these prompts still contained a biased answer - thus demonstrating the propensity of these models to exhibit social bias.
Another interesting finding is that, across all of our experiments (with both models and both decoding strategies), the “doubt bias” prompt style did not have a substantial effect on the proportion of answers containing social bias. For example, for Flan-T5, nucleus sampling, and prompts for which answering “yes” represented bias (Table 7), the original prompt style had 49% of answers containing bias whereas the doubt bias prompt style had 47% . Recall the intention of including the doubt bias prompt style was to test the model’s robustness, since adding statements of uncertainty such as “I’m not sure what to do” should not affect the model’s answer. Hence, we’re able to see that this prompt style displays a level of robustness in the models.
Refer to the Appendix for these results on the rest of the experiments, covering all models and decoding strategies. Note that the trends for both positive bias and doubt bias examples remain the same.
@@
To provide a baseline, we added the “no-stigma” prompt style where we take each of our 37 patterns and remove any mention of a stigma. Hence, we are able to get a sense for whether the models tend to favor the “yes” or “no” answer. On this note, we discover dramatically different proportions when using greedy decoding versus nucleus sampling (Tables 8 and 9). Specifically, for Flan-T5, we discovered that when answering “yes” indicates bias, the model outputted “yes” exactly 0 times during greedy decoding but an average of 36% of the time using nucleus sampling with CoT. Similarly, when answering “no” indicated bias, the model outputted “no” around 22% of the time under greedy decoding and only 12% of the time for nucleus sampling with CoT. Even more interestingly, the results for Flan-UL2 are also different. For both “yes” and “no” as biased answers, the inclusion of nucleus sampling and CoT increases the proportion of biased output when compared with greedy decoding. Refer to the Appendix for further details on results from Flan-UL2.

### 27 Dec 2023 v4

@@ Abstract
Current datasets for unwanted social bias auditing are limited to studying protected demographic features such as race and gender. In this work, we introduce a comprehensive benchmark that is meant to capture the amplification of social bias, via stigmas, in generative language models. Taking inspiration from social science research, we start with a documented list of 93 US-centric stigmas and curate a question-answering (QA) dataset which involves simple social situations. Our benchmark, SocialStigmaQA, contains roughly 10K prompts, with a variety of prompt styles, carefully constructed to systematically test for both social bias and model robustness. We present results for SocialStigmaQA with two open source generative language models and we find that the proportion of socially biased output ranges from 45% to 59% across a variety of decoding strategies and prompting styles. We demonstrate that the deliberate design of the templates in our benchmark (e.g., adding biasing text to the prompt or using different verbs that change the answer that indicates bias) impacts the model tendencies to generate socially biased output. Additionally, through manual evaluation, we discover problematic patterns in the generated chain-of-thought output that range from subtle bias to lack of reasoning.
Warning: This paper contains examples of text which are toxic, biased, and potentially harmful.
@@ 1 Introduction
The rapid advancement of generative language models and their deployment (OpenAI 2022; Bommasani 2023; Nayak 2019; Perspective API 2021) has led to a higher level of scrutiny of their behavior regarding unwanted social bias. Numerous studies have demonstrated that generative models exhibit undesirable behavior that amplifies social bias. Most such studies focus on protected demographic attributes such as gender or race (Baldini et al. 2022; Blodgett et al. 2020; Parrish et al. 2022a; Akyürek et al. 2022). In this work, we go beyond protected demographic attributes and introduce a large dataset that focuses on a comprehensive list of documented US-centric stigmatized conditions (Pachankis et al. 2018a). Examples of stigmatized conditions include being voluntarily childless, living in a trailer park, having facial scars, and more. The dataset is formulated as a question answering task in which questions refer to engaging with a person with a stigmatized condition in a context of a simple social situation. The goal of the question is to probe whether the model generates text that expresses bias against the person with the stigmatized condition. Our experiments with two large open generative models (Wei et al. 2022a; Tay et al. 2023) show that 45%-59% of the questions, depending on the prompting and decoding used, are answered unfavorably towards the person with the stigmatized condition. This result showcases the risk of generative models amplifying biases against people with stigmatized conditions, leading to real harm, as explored by Pachankis et al. (2018a); Mei, Fereidooni, and Caliskan (2023).
Table 1 shows an example from our dataset. All questions are templated such that they can be filled in by the different stigmas considered in this study. The question is preceded by a short context that refers to a simple social situation (e.g., job interview). The question usually asks whether to engage within the social situation with a person that has a stigmatized condition (e.g., hiring someone). To further analyze model behavior, we include two versions for the context: one which expresses an emphasis on how appropriate the person is for the situation (positive bias, e.g., “the candidate who is perfect for the position”) and one in which the person asking the question expresses some doubt or hesitation towards engaging with the person (doubt bias, e.g., “I don’t know what to do.”). Both these versions are meant to investigate if the bias in the generated text diminishes, increases or remains unchanged with the modified context. To establish a baseline, we include a no-stigma version for each template that does not refer to any stigma. These prompts are useful to establish whether certain models have any propensity of answering one way or the other.
@@
    Analysis of chain-of-thought output: Through manual evaluation, we provide high level trends in the generated chain-of-thought output. Specifically, we shed light into a model’s (lack of) reasoning capabilities, which has the potential to exacerbate existing societal inequities by producing biased text in the CoT responses.
@@ 2.1 Social Bias and Stigmas
Social bias can be defined as discrimination for, or against, a person or group, or a set of ideas or beliefs, in a way that is prejudicial or unfair (Webster et al. 2022; Bommasani and Liang 2022). Pachankis et al. (2018b) list 93 different stigmas [^1] An extended version of our paper can be found on arxiv (Nagireddy et al. 2023). The extended version contains the list of all 93 stigmas, additional results and sample chain-of-thought annotations., whilst also documenting the impact of stigmas on health. They consider the definition of stigma as any devalued attribute or characteristic that aims to reduce a person from a whole person to a tainted or discounted one in a particular social context. Stigma affects a substantial segment of the U.S. population at some point of their lives and encompasses a wide range of highly prevalent personal attributes (e.g., old age, obesity, depression) as well as identities or health conditions (e.g., minority sexual orientation, physical disabilities, chronic illnesses). Notably, some stigmas are visible (e.g., facial scars), while others are invisible (e.g., being voluntarily childless).
@@
There is significant work on bias evaluation of language models, such as auditing for unwanted social bias through benchmarks (Baldini et al. 2022; Blodgett et al. 2020; Parrish et al. 2022a; Akyürek et al. 2022; Smith et al. 2022; Selvam et al. 2023; Dhamala et al. 2021; Nangia et al. 2020; Nadeem, Bethke, and Reddy 2020; Wang, Wang, and Yang 2022). Recent efforts propose a holistic evaluation of LMs (Srivastava et al. 2022; Liang et al. 2022) across many datasets, tasks, and metrics. Raji et al. (2021) document the pitfalls of generalizing model ability through a set of benchmarks, while Bowman (2022) discusses the dangers of under-claiming LM abilities. Researchers scrutinized deficiencies of current datasets (Blodgett et al. 2021) and the lack of clarity on the definition of social bias in NLP models and its measures (Blodgett et al. 2020; Selvam et al. 2022). BBQ (Parrish et al. 2022b) is a bias benchmark for QA which utilizes nine social dimensions defined by the US Equal Employment Opportunities Commission (e.g., age, gender identity, physical appearance, etc.). UnQover (Li et al. 2020b) is also a QA dataset that focuses on ambiguous questions for assessing bias across dimensions such as religion, nationality and gender. Smith et al. (2022) introduces a holistic dataset, utilizing a dozen social demographic axes. Importantly, our benchmark offers both a wider variety of categorizations which pertain to stigmas (e.g. voluntarily childless, sex worker, etc.) as well as going deeper into existing segments of these dimensions (e.g. for physical appearance - having limb scars or multiple tattoos, for perceived social status - living in a trailer park, being a gang member, etc.).
We acknowledge that Mei, Fereidooni, and Caliskan (2023) also utilizes the same list of 93 stigmas and analyzes model behavior from the lens of sentiment classification with masked language models. Our work offers two notable differences. First, our benchmark is designed as a question answering task, which enables the evaluation of generative language models in a straightforward way. Second, given the QA task, our benchmark leverages an intentionally constructed assortment of prompts to assess social bias and model robustness to prompting styles. We elaborate on these choices in Section 3.
@@ 3 The SocialStigmaQA Benchmark
The most important structural objects in our benchmark are the 37 patterns present within. These patterns, or templates, were individually hand curated by the team. Each of these patterns follows the same overarching structure - provide a mundane social situation and ask a question pertaining to this situation. Refer to Table 2 for examples of four different patterns. By construction, these patterns contain under-specified contexts (Li et al. 2020a), which means that the ideal answer to each question is some form of “It is impossible to tell without further information.” Note that the biased answer differs depending on the pattern (i.e., it is not always yes or no).
@@
Each member of the team curated patterns in different ways. One member found inspiration for coming up with social situations through existing work on social norms (Forbes et al. 2020). Another team member utilized an experimentation-based strategy by prompting models with different templates in an effort to curate effective patterns. Yet another team member leveraged existing work on discrimination in a variety of settings, including bias in clinical risk prediction (Park et al. 2021; Gilmer and Buccieri 2020), to inform their pattern curation. All members of the team sought to phrase the questions in the patterns such that they could be answered with a yes/no/can’t tell. This was intentional, so as to make automated evaluation possible (refer to Section 4 for more details). There are two important reasons why limiting the answer formulation is preferred. First, an open-ended answer is extremely difficult to analyze automatically, at scale, which does not bode well for a functional benchmark. Second, we are able to get more output than the answer itself by way of chain-of-thought. Hence, by discretizing the answer choices for each question as well as utilizing chain-of-thought, we are able to simultaneously facilitate easy automated evaluation while also retaining longer outputs for manual inspection (more on this in Section 5).
@@
When constructing the patterns, the team diligently worked to reduce any forms of external bias. One example of this was to remove all negative conjunctive adverbs - such as the phrase “however”, which would add an extra layer of doubt that could potentially influence the generated text. To walk through a concrete example, take the following pattern: “My neighbors are in desperate need of help around the house. I was thinking to recommend them my friend who is actually looking for a job. She is someone {stigma}. Shall I recommend them my friend?”. If the pattern was instead phrased as “My neighbors are in desperate need of help around the house. I was thinking to recommend them my friend who is actually looking for a job. However, she is someone {stigma}. Shall I recommend them my friend?,” the inclusion of “However” adds a layer of contradiction, almost equivalent to conditioning the model towards answering in a biased fashion. We aimed to eliminate such characteristics from all the patterns that we created.
@@
An important detail of our benchmark is the intentional attention to two aspects of each template - biased answer choices and prompt styles. First, we carefully construct the benchmark to contain patterns for which answering either “yes” or “no” implies social bias. Refer to Table 2 for concrete examples. Second, we include 4 prompt styles for each pattern:
@@
    doubt bias: the original pattern plus some text that expresses uncertainty. This style aims to test the model’s robustness. In other words, adding text such as “I’m not sure what to do” should not alter the final answer.
@@
Table 1 shows examples of a pattern with the four prompt styles. Refer to the extended version of the paper (Nagireddy et al. 2023) for the full list of 37 patterns, across all four prompt styles.
@@ 4 Experimental Results
We utilized two models - Flan-T5-XXL (11B parameters) (Wei et al. 2022a) and Flan-UL2 (20B parameters) (Tay et al. 2023). We selected these models because they are both open-source, instruction fine-tuned and also trained to produce chain-of-thought output when prompted accordingly. We note that these large language models require GPUs to be hosted and utilized for inference. For each model, we ran greedy decoding as well as nucleus sampling (Holtzman et al. 2019) to produce rich, meaningful CoT. Note that, in general, greedy decoding does not produce meaningful CoT, the generations are usually repetitions of the context. We averaged our results over 5 nucleus sampling runs, each with a different random seed. Across all runs, we obtained 124,320 total answer generations, as follows: 10,360 examples, evaluated with 2 models, and for each model we ran five different seeds for nucleus sampling and one greedy decoding. For prompts run with greedy decoding, we appended “Answer with yes/no/can’t tell” to the end of the prompt. For nucleus sampling, we instead appended “Let’s think step by step” to the end of the prompt to induce chain-of-thought outputs.
@@
A quantitative analysis of the tendencies of the Flan-T5 model to produce a biased answer, by conditioning on both the pattern template format as well as the prompt style, is presented below. We also briefly comment on results with Flan-UL2. Full results for all of our experiments can be found in the extended version of the paper (Nagireddy et al. 2023).
@@
In order to provide the most general quantification of biased output, we report on the proportions of different answers broken down by the types of biased answers. The total proportion of socially biased output on our benchmark ranged from 45% to 59% across a variety of decoding strategies and prompting styles (Table 3). Overall, we noticed that prompting with CoT triggers can either hurt (Flan-T5) or help (Flan-UL2) with the biased answers. Additionally, significant variance was observed across different choices of random seeds, as shown in Figure 1. Bias can vary by more than 10 points across random seeds, a hyperparameter that is usually left to the user of deployed models, or, even worse, randomly set, and, sometimes, not disclosed during deployment - the case of ChatGPT (OpenAI 2022).
Analyzing the results across stigmas, we discovered that the top most biased stigmas belong to categories around sex (e.g., sex offender, having sex for money, genital herpes, HIV) and drug consumption (e.g., drug dealing, cocaine use recreationally). Moreover, stigmas directly referring to race or gender tend to observe less bias. This result emphasizes the importance of expanding model auditing beyond common protected demographic features.
@@
Next, we analyze the impact of triggering chain-of-thought in the generated text by comparing the results from the greedy decoding experiments (Table 4) with results from the nucleus sampling experiments (Table 5). The addition of CoT has mixed results. For the prompts for which answering “no” represents bias, the use of CoT reduces the proportion of such biased answers by around 15% (from 0.692 to 0.535). However, using CoT substantially increases the tendency for bias in prompts for which answering “yes” represents a biased answer by almost 40% (from 0.052 to 0.423). We observe similar trends for Flan-UL2.
@@
Across all of our experiments (with both models and both decoding strategies), the “positive bias” prompt style reduced the proportion of answers containing social bias. This corroborates previous findings in which models produce less harmful and more useful responses when the prompt includes explicit requests to do so (Sun et al. 2023). Compared to original prompt styles, we see that the proportions of biased answers for the positive bias prompt style are smaller. For example, for Flan-T5, greedy decoding, and prompts for which the biased answer is “no” (Table 6), the original prompt style had 84% of answers containing bias whereas the positive bias prompt style had 47%. As mentioned in Section 3, the intention of the positive bias prompt style was specifically to nudge the model towards answering in an unbiased manner - which appears to be the pattern. However, it’s worth noting that even after using this prompt style, approximately half of these prompts still contained a biased answer - thus demonstrating the propensity of these models to exhibit social bias.
On the other hand, across all of our experiments (with both models and both decoding strategies), the “doubt bias” prompt style did not have a substantial effect on the proportion of answers containing social bias. For example, for Flan-T5, nucleus sampling, and prompts for which answering “yes” represented bias (Table 7), the original prompt style had 49% of answers containing bias whereas the doubt bias prompt style had 47% . Recall the intention of including the doubt bias prompt style was to test the model’s robustness, since adding statements of uncertainty such as “I’m not sure what to do” should not affect the model’s answer. Hence, we’re able to see that this prompt style displays a level of robustness in the models.
Our extended version of the paper (Nagireddy et al. 2023) contains results for more experiments, covering all models and decoding strategies. Note that the trends for both positive bias and doubt bias examples remain the same.
@@
To provide a baseline, we added the “no-stigma” prompt style where we take each of our 37 patterns and remove any mention of a stigma. Hence, we are able to get a sense for whether the models tend to favor the “yes” or “no” answer. On this note, we discover dramatically different proportions when using greedy decoding versus nucleus sampling (Tables 8 and 9). Specifically, for Flan-T5, we discovered that when answering “yes” indicates bias, the model outputted “yes” exactly 0 times during greedy decoding but an average of 36% of the time using nucleus sampling with CoT. Similarly, when answering “no” indicated bias, the model outputted “no” around 22% of the time under greedy decoding and only 12% of the time for nucleus sampling with CoT. Even more interestingly, the results for Flan-UL2 are also different. For both “yes” and “no” as biased answers, the inclusion of nucleus sampling and CoT increases the proportion of biased output when compared with greedy decoding.